# Cryogenic geometric anti-spring vibration isolation system

### v1

@@ I Introduction
Scanning probe microscopy (SPM) Battisti et al. (2018) and ultra-sensitive force sensing for quantum gravity experiments Bose et al. (2017) both require ultra-low vibrations at cryogenic temperatures – either to reduce noise in the tunneling current to a few fA Battisti et al. (2018), or to detect effects of gravity with sensitivities in the sub-fN regime Fuchs et al. (2024). The same requirements apply to operational and future interferometer-based gravitational wave detectors, where the mirrors are suspended inside a cryogenic environment Akutsu et al. (2021); Koroveshi et al. (2023); Utina et al. (2022). The challenge of achieving low vibrations is compounded by the growing popularity of closed-system cryostats, driven by helium shortage and correspondingly high prices survey . These systems use pulse tube or GM coolers, which introduce very strong vibrations from pressure changes, typically around $1\text{\,}\mathrm{H}\mathrm{z}$, as can be seen in Fig. 1. Therefore, vibration attenuation especially at low frequencies is the key to realize ultra-sensitive experiments in closed-cycle cryostats.
@@ II.1 Conventional mass-spring systems
The system is characterized by the spring constant $k$ as well as the suspended mass $M$ and spring mass $m$. The spring has an unloaded length $L_{0}$ and we define the loaded spring length $L$ as $L=L_{0}+Mg$, where $g$ is the gravitational acceleration. We further include viscous damping, described by $\gamma$, and structural damping of the spring, described by the loss angle $\phi$. We restrict motion to the vertical $y$-axis and define $y_{0}(t)$ and $y(t)$ as the respective positions of the base and mass w.r.t. some arbitrary origin.
@@
$M\ddot{y}=-k(1+i\phi)\left(L-(y-y_{0})\right)-\alpha m\ddot{y}_{0}-\gamma\dot{y}\text{.}$ (1)
Note that the term $\alpha m\ddot{y}_{0}$ describes the center of percussion effect, which describes the effect of non-aligned forces on both spring ends. As for a simple mass-spring system, $\alpha=0$, this term can be dropped when treating harmonic oscillators. Because it is non-zero in the case of a GAS Stochino et al. (2007), we will include the term here.
@@
The low-pass filtering properties become clear when considering the limit of a massless undamped spring, $m=0,\gamma=0,\phi=0$. At $\omega\ll\omega_{0}$, Eq. 2 reduces to $H_{y}=1$, hence mass and base move in unison. $\omega\approx\omega_{0}$ results in an amplified motion of the mass as $H_{y}$ diverges, and $\omega\gg\omega_{0}$ leads to the desired damping of $\nicefrac{{1}}{{\omega^{2}}}$. Thus, as seen in Eq. 2 the vibration attenuation performance of the mass spring system is optimized by minimizing $\omega_{0}$. For conventional mass-spring systems, reduction of $k$ or increasing of $M$ is cumbersome and competes with spatial constraints, effectively limiting resonance frequencies to around $1\text{\,}\mathrm{H}\mathrm{z}$. For example, a system with $0.5\text{\,}\mathrm{H}\mathrm{z}$ resonance requires a spring extension under load of $1\text{\,}\mathrm{m}$, as one can calculate by equating $F_{g}=Mg$ to $F_{s}=-kx$, and using that $\omega_{0}=\sqrt{\nicefrac{{k}}{{M}}}$.
II.2 Linear GAS model
Next, we will discuss the theory of GAS systems using a linear model, and show that they have the potential to effectively tune the spring constant towards zero, reaching resonance frequencies well below $1\text{\,}\mathrm{H}\mathrm{z}$. We start with a simple two-dimensional model by only considering two compressed blade springs along the horizontal axis, and constraining the movement of the mass in the vertical axis (Fig. 3b). This treatment is based on Blom Blom (2015).
Figure 2: a) To develop the linear model for GAS, the vertical and compressive horizontal force components are separated and represented by conventional springs. The (arbitrary) origin for the vertical coordinates $y$ and $y_{0}$ is indicated at the top of the figure. Adapted from Blom Blom (2015). b) Sketched S-curves in three different regimes. Blue: The undercompressed regime, with a high $k_{\text{eff}}$ in the working point. Green: If the compression is optimally tuned, $k_{\text{eff}}=0$ in the working point, corresponding to the curve being vertical at the working point. Red: In the over-compressed regime, the system is bistable.
In this depiction, the force of each blade can be separated into a horizontal and a vertical component. The linearization of the system is then achieved by modeling these force components as conventional springs as shown in Fig. 2a. In equilibrium, the mass hangs from a vertical spring, and two compressed horizontal springs push radially inward with mutually canceling forces.
Out of equilibrium, the vertical mass-spring system acts to restore the equilibrium position, while the two compressed springs will push the mass further out of equilibrium (See Fig. 2a), hence the name anti-spring. The total vertical force in this configuration is
$F$ $=k_{y}(l_{y}-(y-y_{0}))+k_{x}(l_{x}-l_{x,0})\sin(\theta)$ (3)
$=\left(k_{y}+k_{x}\left(1-\frac{l_{x,0}}{l_{x}}\right)\right)(l_{y}-(y-y_{0}))\text{,}$ (4)
where $k_{y}$ is the spring constant for the vertical spring and $k_{x}$ the combined spring constant of the two individual horizontal springs. In equilibrium, the vertical spring has length $l_{y}$. $l_{x,0}$ denotes the uncompressed length of the horizontal springs and $l_{x}$, $x_{0}$ can be inferred from Fig. 2a.
Based on Eq. 4 we define an effective spring constant which can be rewritten as
$k_{\text{eff}}(y)=k_{y}+k_{x}\left(1-\frac{l_{x,0}}{\sqrt{(l_{y}-(y-y_{0}))^{2}+x_{0}^{2}}}\right)\text{.}$ (5)
From this equation it becomes clear that the spring constant can be modified by tuning $x_{0}$ and thereby the compression of the horizontal springs. The situation of minimal compression is called the working point of the filter. In this configuration, the minimal resonance frequency ensures optimal filtering properties. In principle, even zero stiffness could be achieved, providing ideal filtering behavior.
@@ II.2 Linear GAS model
Eq. 5 also shows the possibility of $k_{\text{eff}}<0$ for some range of $y$ if the springs are compressed beyond the working point. In this situation, the system has three equilibrium positions, the middle one of which is unstable. Under sufficiently large forces, the mass can jump between the top and the bottom position.
@@
Experimentally, we extract the effective spring constant from measurements of the so-called S-curve. This curve shows the equilibrium position as a function of load $M$. The effective spring constant is inversely proportional to the slope of the S-curve. The working point is the point where the S-curve is steepest.
In this model, the transfer function of a GAS filter follows Eq. 2, where we need to consider the non-zero center-of-percussion effect. For a monolithic GAS, it is characterized by $\nicefrac{{1}}{{5}}\leq\alpha\leq\nicefrac{{1}}{{8}}$, depending on blade geometry Stochino et al. (2007). In our system, we have an overall blade mass $m$ of around 60 g, leading to saturation of the $\nicefrac{{1}}{{\omega^{2}}}$-decay for frequencies $\gtrsim$5\text{\,}\mathrm{H}\mathrm{z}$\$. Note that Stochino et al. Stochino et al. (2007) proposed a method to compensate the center-of-percussion effect, lowering the saturation level significantly.
@@ III Design of a cyrogenic GAS and cryostat
For our cryogenic vibration attenuation system, we chose a monolithic GAS design that is well-established in gravitational wave detectors Akutsu et al. (2021); Beker et al. (2012). It consists of quasi-triangular blade springs, arranged in a radial geometry (Fig 3a). The outer end of each blade is anchored in its own base, the inner side of the blades come together and are joined at the key stone, from where the mass is suspended.
An exact 1 dimensional model of this geometry has been worked out by Cella et al. Cella et al. (2005), but the linearized model presented above describes the key characteristics. The main difference between both models are the exact dynamics at the working point. For instance, the ability to reach $k_{\text{eff}}=0$ depends heavily on the blades.
@@ III.1 Blade design
(i) The shape of the blades determines the bending profile under load. As the resonance frequency $\omega_{0}\propto M^{-\nicefrac{{1}}{{2}}}$, it seems natural to design a blade shape that maximizes the load it can carry before exceeding the material’s yield strength. Furthermore, the saturation level of the $\nicefrac{{1}}{{\omega^{2}}}$ decay is decreased by reducing the ration $\nicefrac{{m}}{{M}}$ (See Eq. 2). Load maximization for limited blade size is achieved by the so-called constant-curvature solution Blom (2015); Cella et al. (2005), where the blade bends uniformly when loaded. However, as shown by Cella et al. Cella et al. (2005), this blade shape will not be able to reach the desired zero stiffness. They propose a shape that does reach zero stiffness, while approximating the constant curvature solution.
@@
Our blades are designed based on these three considerations. The blade profile $\Gamma$ follows a fifth-order polynomial as a function of $\nicefrac{{x}}{{L}}$, the normalized coordinate along the blade length $x$. $\Gamma$ is given by:
$\Gamma\left(\frac{x}{L}\right)=W_{0}\sum_{i=0}^{5}c_{i}\left(\frac{x}{L}\right)^{i}\text{,}$ (6)
where the coefficients $c_{0}=0.8317,c_{1}=0.9328,c_{2}=-1.9631,c_{3}=5.8534,c_{4}=-12.22,c_{5}=6.6757$. $W_{0}$ is the width at the spring base. In our system the dimensions are $W_{0}=$33.55\text{\,}\mathrm{m}\mathrm{m}$\$ and $L=$133.64\text{\,}\mathrm{m}\mathrm{m}$\$. These were chosen to maximize the blade length and carrying strength, given the limited space inside the cryostat. The blades are made of $0.6\text{\,}\mathrm{m}\mathrm{m}$ thick grade 5 titanium (Ti-6Al-4V), which has a yield strength exceeding $1\text{\,}\mathrm{G}\mathrm{P}\mathrm{a}$. In their working point, the springs are compressed, such that they depart from their base at an angle of $45\degree$ and reach the key stone at an angle of $33\degree$ with respect to the horizontal plane (Fig. 3b).
@@ IV Performance
Figure 4: An iterative tuning process for our GAS filter based on recording S-curves, position-load diagrams. Initially, the compression is too low, so it is increased by shifting the spring bases inward by $\Delta x_{0}$ until reaching a near-vertical slope. The voice coil current is converted to a mass difference $\Delta M$.
@@ IV.1 Tuning the system
Figure 5: Characterization of the GAS filter. a) The S-curve of the tuned filter. b) For the tuned filter, the resonance frequency is recorded as a function of the applied load. We record a minimal frequency of $185\text{\,}\mathrm{m}\mathrm{H}\mathrm{z}$. c) As a function of temperature, the slope at the working point, shown in the top panel, undergoes slight changes due to the different thermal expansion coefficients of the blades and the base. Much more pronounced is the decreasing working point load due to the reducing Young’s modulus, shown in the bottom panel.

### v2

@@ I Introduction
Scanning probe microscopy (SPM) Battisti et al. (2018) and ultra-sensitive force sensing for quantum gravity experiments Bose et al. (2017) both require ultra-low vibrations at cryogenic temperatures – either to reduce noise in the tunneling current to a few fA Battisti et al. (2018), or to detect effects of gravity with sensitivities in the sub-fN regime Fuchs et al. (2024). The same requirements apply to operational and future interferometer-based gravitational wave detectors, where the mirrors are suspended inside a cryogenic environment Akutsu et al. (2021); Koroveshi et al. (2023); Utina et al. (2022). The challenge of achieving low vibrations is compounded by the growing popularity of closed-system cryostats, driven by helium shortage and correspondingly high prices noa (2025). These systems use pulse tube or GM coolers, which introduce very strong vibrations from pressure changes, typically around $1\text{\,}\mathrm{H}\mathrm{z}$, as can be seen in Fig. 1. Therefore, vibration attenuation especially at low frequencies is the key to realize ultra-sensitive experiments in closed-cycle cryostats.
@@ II.1 Conventional mass-spring systems
The system is characterized by the spring constant $k$ as well as the suspended mass $M$ and spring mass $m\ll M$. The spring has a loaded length $l$. We further include viscous damping, described by $\gamma$, and structural damping of the spring, described by the loss angle $\phi$. We restrict motion to the vertical $y$-axis and define $y_{0}(t)$ and $y(t)$ as the respective vertical coordinates of the base and mass w.r.t. some inertial frame.
@@
$M\ddot{y}=-k(1+i\phi)\left((y-(y_{0}-l)\right)-\alpha m\ddot{y}_{0}-\gamma\dot{y}\text{.}$ (1)
The term $\alpha m\ddot{y}_{0}$ describes the center-of-percussion effect, arising in case of non-aligned forces on both spring ends. In the case of a coil spring, this term vanishes as $\alpha=0$. Because it is non-zero in the case of a GAS Stochino et al. (2007), we will include the term here.
@@
The low-pass filtering properties become clear when considering the limit of a massless undamped spring, $m=0,\gamma=0,\phi=0$. At $\omega\ll\omega_{0}$, Eq. 2 reduces to $H_{y}=1$, hence mass and base move in unison. $\omega\approx\omega_{0}$ results in an amplified motion of the mass as $H_{y}$ diverges, and $\omega\gg\omega_{0}$ leads to the desired damping of $\nicefrac{{1}}{{\omega^{2}}}$. Thus, as seen in Eq. 2 the vibration attenuation performance of the mass spring system is optimized by minimizing $\omega_{0}$. For conventional mass-spring systems, reduction of $k$ or increasing of $M$ is cumbersome and competes with spatial constraints, effectively limiting resonance frequencies to around $1\text{\,}\mathrm{H}\mathrm{z}$. For example, a linear mass-spring system with a resonance frequency of 0.5 Hz shows a spring extension under load of $\nicefrac{{Mg}}{{k}}=\nicefrac{{\omega_{0}^{2}}}{{g}}\approx$1\text{\,}\mathrm{m}$\$. A GAS spring, tuned to such a frequency occupies only a small fraction of vertical space, thanks to its non-linear spring behavior
II.2 Geometric anti-spring systems
Next, we will discuss the theory of GAS systems, by constructing a model of our GAS linear springs. We show that their effective spring constant can be tuned to extremely low values, reaching resonance frequencies well below 1 Hz.
Our GAS uses compressed blade springs placed radially symmetric around a key stone, from which a mass $M$ can be suspended (Fig. 3a). Using the radial symmetry, we can characterize the system in two dimensions by considering a cross section of two opposing blades, as indicated in Fig. 3b. In the following, we treat the horizontal and vertical force components separately and constrain movement to the vertical direction.
In the model of Fig. 2a, the vertical force component is obtained from a linear coil spring with stiffness $k_{y}$, carrying the payload. The compressive horizontal component $F_{c}$ acts on the keystone through symmetrically configured linear springs. We define the working height as the situation in which the compressive forces act purely horizontally on the key stone, canceling each other out. In this model, this corresponds to the compressed springs being horizontal.
Suppose the system is at rest with the key stone at the working height and is subject to a displacement $\delta_{y}$ from working height (See Fig. 2a). The vertical spring counteracts the deviation, while the compressed springs do the opposite, pushing the key stone further from the working point, hence the name anti-spring. The net vertical force $\delta F_{y}=F_{y}-Mg$ in this configuration is
$\delta F_{y}$ $=-k_{y}\delta_{y}+NF_{c}\sin(\theta)$ (3)
$=-\left(k_{y}-\frac{NF_{c}}{\sqrt{L_{x}^{2}+\delta_{y}^{2}}}\right)\delta_{y}\text{,}$ (4)
where $k_{y}$ is the spring constant for the vertical spring and $N$ refers to the number of blade springs in the GAS. In this two-dimensional model, N=2{N=2}. The angle $\theta$ and distance $L_{x}$ can be inferred from Fig. 2a.
Based on Eq. 3 we approximate the effective spring constant for $\delta_{y}\ll L_{x}$ by differentiating $\delta F_{y}$ w.r.t. $\delta_{y}$
$k_{\text{eff}}(\delta_{y})\approx k_{y}-\frac{NF_{c}}{\sqrt{L_{x}^{2}+\delta_{y}^{2}}}\text{.}$ (5)
Figure 2: a) To develop the linear model for GAS, the vertical and compressive horizontal force components are separated. The vertical force is represented as a simple coil spring. The horizontal forces are transferred to the key stone via compressed springs. At the working height, the compressed springs are horizontal and the compressive forces cancel each other out. Adapted from Blom Blom (2015). b) S-curves following Eq. 5 for different compressive forces $F_{c}$. At $\delta y=0$, all curves obtain minimal effective stiffness $k_{\text{eff}}$. The critical green curve has exactly $k_{\text{eff}}(\delta y)=0$. The red curve represents the bistable state, where the minimal stiffness is negative.
From this equation it becomes clear that the spring constant can be modified by tuning the compressive force $F_{c}$.
The goal of GAS filters is to tune them to their working point, where $F_{c}$ is chosen so that $k_{\text{eff}}$ is sufficiently low, but positive around the working height $\delta_{y}=0$. This configuration ensures optimal filtering performance.
@@ II.2 Geometric anti-spring systems
If the springs are compressed beyond their working point, $k_{\text{eff}}<0$ for some range around $\delta_{y}=0$. Within this range, there are two equilibrium positions symmetrically around $\delta_{y}=0$. Under sufficiently large forces, the mass can jump between the top and the bottom position.
@@
Experimentally, we extract the effective spring constant from measurements of the so-called S-curve. This curve shows the equilibrium position as a function of load $M$. The effective spring constant is inversely proportional to the slope of the S-curve, having its minimum at $\delta_{y}=0$. Fig. 2b shows theoretical curves for three different values of $F_{c}$, where in the bistable regime, the curve is multi-valued.
The stiffness $k_{\text{eff}}$ can be used to describe the GAS transfer function using Eq. 2. For the GAS, the center-of-percussion effect leads to saturation of the $\nicefrac{{1}}{{\omega^{2}}}$-attenuation. $\alpha$ characterizes the strength of the effect, and is dependent on the blade geometry. Stochino et al. Stochino et al. (2007) proposed a method to compensate the center-of-percussion effect, lowering the saturation level significantly.
@@ III Design of a cyrogenic GAS and cryostat
For our cryogenic vibration attenuation system, we chose a monolithic GAS design that is well-established in gravitational wave detectors Akutsu et al. (2021); Beker et al. (2012). It consists of quasi-triangular blade springs, arranged in a radial geometry (Fig 3a), clamped between a base and the key stone, from where the mass is suspended.
An exact 1 dimensional model of this geometry has been worked out by Cella et al. Cella et al. (2005), but the linearized model presented above describes the key characteristics.
@@ III.1 Blade design
(i) The shape of the blades determines the bending profile under load. As the resonance frequency $\omega_{0}\propto M^{-\nicefrac{{1}}{{2}}}$, it seems natural to design a blade shape that maximizes the load it can carry before exceeding the material’s yield strength. Furthermore, the saturation level of the $\nicefrac{{1}}{{\omega^{2}}}$ decay is decreased by reducing the ratio $\nicefrac{{m}}{{M}}$ (See Eq. 2). Load maximization for limited blade size is achieved by the so-called constant-curvature solution Blom (2015); Cella et al. (2005), where the blade bends uniformly when loaded. However, as shown by Cella et al. Cella et al. (2005), this blade shape will not be able to reach the desired zero stiffness. They propose a shape that does reach zero stiffness, while approximating the constant curvature solution.
@@
Our blades are designed based on these three considerations. The blade profile $W$ follows a fifth-order polynomial as a function of $\nicefrac{{s}}{{L}}$, where $0\leq s\leq L$ is the curvilinear coordinate along the blade. $W$ is given by:
$W\left(\frac{s}{L}\right)=W_{\text{ref}}\sum_{i=0}^{5}c_{i}\left(\frac{s}{L}\right)^{i}\text{,}$ (6)
where the coefficients $c_{0}=0.8317,c_{1}=0.9328,c_{2}=-1.9631,c_{3}=5.8534,c_{4}=-12.22,c_{5}=6.6757$. $W_{\text{ref}}$ characterizes the width of the blade. In our system the dimensions are $W_{\text{ref}}=$33.55\text{\,}\mathrm{m}\mathrm{m}$\$ and $L=$133.64\text{\,}\mathrm{m}\mathrm{m}$\$. These were chosen to maximize the blade length and carrying strength, given the limited space inside the cryostat. The blades are made of $0.6\text{\,}\mathrm{m}\mathrm{m}$ thick grade 5 titanium (Ti-6Al-4V), which has a yield strength exceeding $1\text{\,}\mathrm{G}\mathrm{P}\mathrm{a}$. In their working point, the springs are compressed, such that they depart from their base at an angle of $45^{\circ}$ and reach the key stone at an angle of $33^{\circ}$ with respect to the horizontal plane (Fig. 3b).
@@ IV Performance
Figure 4: An iterative tuning process for our GAS filter based on recording S-curves, position-load diagrams. Initially, the compression is too low, so it is increased by shifting the spring bases inward by $\Delta L_{x}$ until reaching a near-vertical slope. The voice coil current is converted to a mass difference $\Delta M$.
@@ IV.1 Tuning the system
Figure 5: Characterization of the GAS filter. a) The S-curve of the tuned filter. Note that the zero-point $y=0$ is chosen differently from the data presented in Fig. 4. b) For the tuned filter, the resonance frequency is recorded as a function of the applied load. We record a minimal frequency of $185\text{\,}\mathrm{m}\mathrm{H}\mathrm{z}$. c) As a function of temperature, the slope at the working point, shown in the top panel, undergoes slight changes due to the different thermal expansion coefficients of the blades and the base. Much more pronounced is the decreasing working point load due to the reducing Young’s modulus, shown in the bottom panel.
@@ V Conclusions
Acknowledgements
This work was supported by the European Research Council Consilidator Grant (Project PairNoise).
Author declarations
Conflicts of Interest
The authors have no conflicts to disclose.
Author contributions
L. Feenstra: Conceptualization (equal); Data curation (equal); Formal analysis (equal); Investigation (equal); Methodology (equal); Project administration (equal); Software (equal); Supervision (equal); Validation (equal); Visualization (equal); Writing – original draft (equal); Writing – review & editing. S. Domínguez-Calderón: Data curation (equal); Formal analysis (equal); Investigation (equal); Methodology (equal); Software (equal); Validation (equal); Writing – review & editing. K. van Oosten: Investigation (equal); Methodology (equal); Writing – review & editing. H.S.M. Bohemen: Investigation (equal); Methodology (equal); Writing – review & editing. T. Benschop: Conceptualization (equal); Formal analysis (equal); Investigation (equal); Methodology (equal); Supervision (equal); Validation. M. Brinkman: Investigation (equal); Methodology (equal); Project administration (equal); Validation. M. Li: Data curation (equal); Formal analysis (equal); Investigation (equal); Methodology (equal); Project administration (equal); Supervision (equal); Validation. E. Hennes: Conceptualization (equal); Methodology (equal); Writing – review & editing. R. Cornelissen: Conceptualization (equal); Methodology (equal); Writing – review & editing. B.J. Hensen: Conceptualization (equal); Funding acquisition (equal); Methodology (equal); Project administration (equal); Supervision (equal); Validation (equal); Writing – review & editing. A. Bertolini: Conceptualization (equal); Methodology (equal); Validation (equal); Writing – review & editing. M.P. Allan: Conceptualization (equal); Funding acquisition (equal); Methodology (equal); Project administration (equal); Supervision (equal); Validation (equal); Writing – original draft (equal); Writing – review & editing.
Data Availability Statement
The data that support the findings of this study are openly available in [tbd], at [tbd], reference number [tbd].